Find the general antiderivative of 3 + 5x +10 (3) b) Sinx + 5sex
Answer #1

Que 3 3 Aus! Autideivative > gutegrat roction +5x +io)dx x 4 + +10%.+C. 3/2 Y2+1 a +1oxte [ f(x) = x + 3 Aus- b) f(x) = {(sin

If you have any questions please let me know

Thank You

Please give me thumb up......